OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a twelve-inch diameter pizza and an eighteen-inch diameter pizza. He has set the price of a twelve-inch cheese pizza at eight dollars, based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an eighteen-inch cheese pizza?

OpenStudy (anonymous):

@JA1 HELP?

OpenStudy (anonymous):

($8)/(12 inches) = ($x)/(18 inches) You can (should) eliminate the $ and "inches": 8/12 = x/18 So, just cross-multiply and solve for "x". Can you do that or do you need help with that?

OpenStudy (anonymous):

This format might make it easier to see:\[\frac{ 8 }{ 12 } = \frac{ x }{ 18 }\]

OpenStudy (anonymous):

(12)(x) = (8)(18) 12x = 144 -> x = 12 (dollars)

OpenStudy (anonymous):

All good now, @Ryleighblue ?

OpenStudy (anonymous):

Hold on hold on hold on I'm very confuzzled let me focus

OpenStudy (anonymous):

I have no clue what my answer is doe

OpenStudy (anonymous):

12 uw! Good luck to you in all of your studies and thx for the recognition! @Ryleighblue

OpenStudy (anonymous):

That's because : 8/12 = 12/18 They both equal 2/3